# A fast particle-field method for 3D Chemotaxis systemsJingyuan Hu, Zhongjian Wang, Jack Xin and Zhiwen Zhang

### v2

@@ Abstract
In this paper, we develop a novel numerical framework, namely the stochastic interacting particle-field method with particle-in-cell acceleration (SIPF-PIC), for the efficient simulation of the three-dimensional (3D) parabolic-parabolic Keller-Segel (KS) systems. The SIPF-PIC method integrates Lagrangian particle dynamics with spectral field solvers by leveraging localized particle-grid interpolations and fast Fourier transform (FFT) techniques. For $P$ particles and $H$ Fourier modes per spatial dimension, the SIPF-PIC method achieves a computational complexity of $O(P+H^{3}\log H)$ per time step, a significant improvement over the original SIPF method (proposed in [SIPF1]), which has a computational complexity of $O(PH^{3})$, while preserving numerical accuracy. Moreover, we carry out a rigorous error analysis for the proposed method and establish the corresponding error estimates. Finally, we present numerical experiments to validate the convergence order and demonstrate the computational efficiency of SIPF-PIC. Further numerical experiments show the method’s capability of capturing complex blowup dynamics beyond single-point collapse, including ring-shaped singularities.
@@ 1 Introduction
Due to the nonlinear and potentially singular behavior of KS systems, especially in the presence of blow-up phenomena [nagai1995blow, herrero1998self], numerical methods have become essential for analyzing their solutions. Mesh-based numerical methods are among the most widely used approaches for solving the KS system. Chertock et al. [pos_preserving] proposed a high-order hybrid finite volume and finite difference scheme with positivity-preserving properties for two-dimensional (2D) problems. Shen et al. [shen2020unconditionally] proposed an energy dissipation and bound-preserving scheme that is not restricted to specific spatial discretizations. Chen et al. [chen2022error] developed a fully-discrete finite element scheme for the 2D parabolic-elliptic case, and showed that the scheme exhibits finite-time blow-up under assumptions analogous to the continuous case. Other works include a positivity preserving and asymptotic preserving semi-discrete scheme [LiuJian-Guo2018Paap], and a saturation concentration setting [hillen2001global] which prevents blow-up and exhibits spiky solutions. Beyond these 2D cases, Gnanasekaran et al. [3Dblowup] simulated a 3D blow up solution in a cube with a finite difference scheme in a short time.
@@
Existing Lagrangian frameworks for parabolic-elliptic cases cannot be directly extended to fully parabolic systems, as the chemo-attractant concentration no longer achieves rapid local equilibration and the drift terms depend on historical particle positions [chen2022mean, fournier2023particle]. Stevens [stevens2000] developed an $N$-particle system for the fully parabolic case, representing the chemo-attractant concentration with particles subject to probabilistic creation and annihilation. However, this method also leads to the accumulation of per-step complexity over time. To address this limitation, Wang et al. [SIPF1] developed a stochastic interacting particle-field (SIPF) method that records the chemo-attractant concentration through spectral decomposition and maintains only the current-step particle positions, ensuring that the complexity per step remains constant over time.
@@
However, we have to point out that in the SIPF method [SIPF1], each particle of $\rho$ interacts with every Fourier mode of $c$. This requires the computation of the trigonometric sum $\frac{1}{P}\sum_{p=1}^{P}e^{-\frac{2\pi i\mathbf{q}}{L}\cdot\widehat{\mathbf{X}}_{p}}$ for all wavenumbers $\mathbf{q}$, where $\widehat{\mathbf{X}}_{p}$ for $1\leq p\leq P$ are the positions of the particles (see Algorithm 1). For 3D systems with $H$ Fourier modes per dimension, this leads to an $O(PH^{3})$ computational complexity per timestep. Such a computational bottleneck severely restricts the applicability of SIPF for large-scale 3D KS systems, particularly in scenarios requiring high-resolution and long-time simulations.
@@
To reduce the complexity, in this paper, we propose the SIPF-PIC framework (see Algorithm 4) that integrates localized particle-grid operations with spectral field solvers accelerated by the fast Fourier transform (FFT). This hybrid method confines particle interactions to adjacent grid nodes while resolving global frequency-domain couplings via FFT, yielding a computational complexity of $O\left(P+H^{3}\log H\right)$ per timestep, as presented in Theorem 3.7. The bidirectional interaction mechanism operates in each time step through two stages: Particle-to-Grid projection (see Algorithm 2) transfers particle-derived data to the computational grid, while Grid-to-Particle interpolation (see Algorithm 3) distributes field solutions from the grid back to individual particles; details are provided in Section 3. The Particle-to-Grid projection in the SIPF-PIC framework, which numerically realizes the trigonometric summation that is also known as the unequally spaced fast Fourier transform (USFFT) [USFFT], has been employed in Fourier-basis particle simulations [USFFTS]. The Grid-to-Particle interpolation is employed in classical particle-in-cell methods [alma99, TskhakayaD.2007TPM] and does not rely on spectral decompositions. Our algorithm reduces the particle-grid interaction complexity from $O\left(PH^{3}\right)$ to $O\left(P+H^{3}\log H\right)$, enabling the simulation of 3D KS systems with significantly increased particle numbers $(P>10^{6})$ and finer grids $(H=256)$, and thus resulting in improved numerical accuracy.
@@
We also present numerical experiments to verify the convergence rate, computational complexity, and the consistency with respect to $H$. We verify the critical mass $8\pi$ and the evolution of the second moment in a standard two-dimensional parabolic-elliptic system (1), where the dimension $d=2$ and the parameters $(k,\epsilon,\mu,\chi)=(0,0,1,1)$. For 3D parabolic-parabolic KS systems with $k,\epsilon>0)$, the critical mass lacks a precise characterization due to its dependence on the initial density distribution $\rho(\mathbf{x},0)$. We validate the accuracy of SIPF-PIC under radially symmetric initial data, where the reference solution can be obtained by reducing the system to 1D and running a finite difference method on a fine mesh.
@@
Moreover, the SIPF-PIC successfully reproduces complex blow-up patterns beyond the classical single-point radial type blow-up. We first investigate a non-radially symmetric configuration with four uniform density balls located at the vertices of a tetrahedron, and then accurately resolve the ring-shaped blow-up pattern proposed in a recent paper [ring_ref], where mass concentrates on an evolving ring $x_{1}^{2}+x_{2}^{2}=r(t)^{2}$ with a time-dependent radius $r(t)$. In contrast, the original SIPF, when using the parameter scaling ($P=10^{4},H=24$), does not have enough resolution to capture the ring-shaped blow-up pattern. Specifically, the SIPF-PIC successfully observes the line-density blow-up, which is characterized by mass concentration along evolving spatial structures rather than isolated points, and further captures its instability under symmetry-breaking perturbations [collapsing]. These complex collapse mechanisms deviate fundamentally from classical delta-type blow-up dynamics, and the SIPF-PIC, by leveraging large particle numbers and high-resolution grids, reliably captures their evolutionary process in simulations.
@@
The rest of the paper is organized as follows. Section 2 reviews theoretical finite-time blow-up results for the KS system (1), and introduces the SIPF, a hybrid particle-spectral method derived through temporal discretization of the system. In Section 3, we formalize the interpolation operators governing bidirectional particle-grid coupling, including Particle-to-Grid discretization (commonly termed USFFT) and Grid-to-Particle interpolation, followed by the enhanced SIPF-PIC method. Subsection 4 rigorously quantifies a comprehensive error analysis framework that rigorously quantifies cumulative approximation errors from various sources in the SIPF-PIC method. Section 5 validates the method through numerical experiments, including convergence tests and performance benchmarks, as well as the visualization of 3D blow-up patterns.
@@ 2.1 Analytical results of KS system
(3) $\frac{d}{dt}m_{2}(t)=4M_{0}\left(1-\frac{M_{0}}{8\pi}\right),$
@@ 2.2 SIPF Method
Throughout this section, the system is restricted to a domain $\Omega=\left[-\frac{L}{2},\frac{L}{2}\right]^{3}$ with periodic boundary conditions. In the SIPF method for solving the parabolic-parabolic KS system (2), the chemical concentration $\hat{c}(\mathbf{x},n\tau)$ is represented in frequency space using a Fourier series expansion, while the particle density $\hat{\rho}(\mathbf{x},n\tau)$ is represented in physical space by some moving particles. The temporal domain $[0,T]$ is partitioned into $N_{T}=\frac{T}{\tau}$ equal intervals of size $\tau$.
@@
(5) $\hat{c}^{(n)}(\mathbf{x}):=\hat{c}(\mathbf{x},n\tau)=\sum_{\mathbf{q}\in U}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}},$
@@
where $\delta(\cdot)$ is the Dirac delta function, $M_{0}=\int_{\Omega}\rho^{(0)}(\mathbf{x})\,\mathrm{d}\mathbf{x}$ is the initial mass, and $(\widehat{\mathbf{X}}_{p}^{(n)})_{1\leq p\leq P}$ are evolving particles.
@@
(9) $\hat{\beta}_{\mathbf{q}}^{(n-1)}$ $=\frac{1}{L^{3}}\int_{\Omega}\hat{\rho}^{(n-1)}(\mathbf{x})e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}\,\mathrm{d}\mathbf{x}$
$=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}e^{-\frac{2\pi i}{L}\mathbf{q}\cdot\widehat{\mathbf{X}}_{p}^{(n-1)}}.$
@@
(11) $\mathrm{d}\mathbf{X}=\chi\nabla_{\mathbf{X}}c(\mathbf{X},t)\mathrm{d}t+\mathrm{d}\mathbf{W}_{t}$
@@
where $\mathbf{W}_{p}^{(n)}\sim\mathcal{N}(0,\,2\mu\tau I_{3})$ are independent 3D Wiener increments, and $I_{3}$ is the $3\times 3$ identity matrix, and the drift $\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})$ is computed by
@@ 3.1 Particle-in-Cell Interpolation Framework
The main drawback of the original SIPF (Algorithm 1) is the particle-frequency interaction with $O(PH^{3})$ computational complexity. Our new framework is designed to break through this computational bottleneck.
@@
Consider a smooth target function $f:\mathbb{R}^{3}\to\mathbb{R}$. Let $\overline{\mathbf{p}}$ be any three-dimensional real index. Its corresponding position in physical space is $\overline{\mathbf{p}}\frac{L}{H}$, that is, we use the notation $f_{\overline{\mathbf{p}}}=f_{\overline{p}_{1},\overline{p}_{2},\overline{p}_{3}}:=f\left(\overline{\mathbf{p}}\frac{L}{H}\right)$. Let $\mathbf{p}$ be the closest three-dimensional integer index to $\overline{\mathbf{p}}$, with each component less than the corresponding component of $\overline{\mathbf{p}}$. Specifically, for $\overline{\mathbf{p}}=(\overline{p}_{1},\overline{p}_{2},\overline{p}_{3})=\mathbf{p}+\boldsymbol{\lambda}=(p_{1}+\lambda_{1},p_{2}+\lambda_{2},p_{3}+\lambda_{3})$ where $\mathbf{p}\in\mathbb{Z}^{3}$ and $0\leq\lambda_{1},\lambda_{2},\lambda_{3}\leq 1$, a conventional particle-in-cell (PIC) method [alma99] represents $f_{\overline{\mathbf{p}}}$ as a convex combination of the eight vertices of the cell it occupies for interpolation:
(14) $f_{\overline{\mathbf{p}}}\approx$ $\ (1-\lambda_{1})(1-\lambda_{2})(1-\lambda_{3})f_{p_{1},p_{2},p_{3}}+(1-\lambda_{1})(1-\lambda_{2})\lambda_{3}f_{p_{1},p_{2},p_{3}+1}$
$\ +(1-\lambda_{1})\lambda_{2}(1-\lambda_{3})f_{p_{1},p_{2}+1,p_{3}}+(1-\lambda_{1})\lambda_{2}\lambda_{3}f_{p_{1},p_{2}+1,p_{3}+1}$
$\ +\lambda_{1}(1-\lambda_{2})(1-\lambda_{3})f_{p_{1}+1,p_{2},p_{3}}+\lambda_{1}(1-\lambda_{2})\lambda_{3}f_{p_{1}+1,p_{2},p_{3}+1}$
$\ +\lambda_{1}\lambda_{2}(1-\lambda_{3})f_{p_{1}+1,p_{2}+1,p_{3}}+\lambda_{1}\lambda_{2}\lambda_{3}f_{p_{1}+1,p_{2}+1,p_{3}+1}.$
For any position $\mathbf{X}=\overline{\mathbf{p}}\frac{L}{H}$, to simplify the expression of interpolation, we denote:
(15) $\mathbf{X}_{\boldsymbol{\gamma}}=\left(\left\lfloor\frac{H}{L}\mathbf{X}\right\rfloor+\boldsymbol{\gamma}\right)\times\frac{L}{H},$
@@
where the notation $\lfloor\cdot\rfloor$ denotes the floor function applied component-wise. Define the second-order basis $\Gamma_{2}=\{0,1\}^{3}$. The interpolation (14) can be compactly expressed as
(16) $f(\mathbf{X})\approx\sum_{\boldsymbol{\gamma}\in\Gamma_{2}}F_{\boldsymbol{\gamma}}^{(2)}f(\mathbf{X}_{\boldsymbol{\gamma}})=F_{000}^{(2)}f(\mathbf{X}_{000})+\cdots+F_{111}^{(2)}f(\mathbf{X}_{111}),$
where the weights $F_{j_{1}j_{2}j_{3}}^{(2)}=\prod_{s=1}^{3}\left(1-j_{s}+(2j_{s}-1)\lambda_{s}\right)$ for $(j_{1},j_{2},j_{3})\in\{0,1\}^{3}$ correspond to (14). This trilinear interpolation exhibits second-order truncation error, and the eight vertices $\mathbf{X}_{000},\ldots,\mathbf{X}_{111}$ exhibit cubic symmetry.
@@
For points near the boundary under periodic boundary conditions, the indices are wrapped around the grid edges.
Figure 1: Extended vertices for fourth-order interpolation
To achieve higher precision, we also introduce 24 new grid points $\Gamma_{4}$ adjacent to the original 8 grid points, where two of $\gamma_{1},\gamma_{2},\gamma_{3}$ are 0 or 1, and the remaining one is -1 or 2, as shown in Figure 1. These extended nodes enable interpolation [tricubic] exhibiting fourth-order truncation error. Vertices $X_{\gamma}$ are extended to $\gamma\in\Gamma_{4}$, and their coefficients $F_{\gamma}^{(4)}$ are shown below, exhibiting cubic symmetry:
(17) $F_{2,j_{2},j_{3}}^{(4)}=-\frac{(1+\lambda_{1})(1-\lambda_{1})\lambda_{1}\left(1-j_{2}+(2j_{2}-1)\lambda_{2}\right)\left(1-j_{3}+(2j_{3}-1)\lambda_{3}\right)}{6},$
$F_{-1,j_{2},j_{3}}^{(4)}=-\frac{(2-\lambda_{1})\lambda_{1}(1-\lambda_{1})\left(1-j_{2}+(2j_{2}-1)\lambda_{2}\right)\left(1-j_{3}+(2j_{3}-1)\lambda_{3}\right)}{6},\cdots$
(18) $F_{j_{1}j_{2}j_{3}}^{(4)}=\prod_{s=1}^{3}\left(1-j_{s}+(2j_{s}-1)\lambda_{s}\right)\left(1+\sum_{k=1}^{3}\frac{\lambda_{k}(1-\lambda_{k})}{2}\right),(j_{1},j_{2},j_{3})\in\{0,1\}^{3}$
The complete fourth-order interpolation formula becomes:
(19) $f(\mathbf{X})\approx\sum_{\gamma\in\Gamma_{4}}F_{\gamma}^{(4)}f(\mathbf{X}_{\gamma}),$
where
$\Gamma_{4}=\Gamma_{2}\cup\{-1,2\}\times\{0,1\}\times\{0,1\}\cup\{0,1\}\times\{-1,2\}\times\{0,1\}\cup\{0,1\}\times\{0,1\}\times\{-1,2\}.$
@@
Let $\mathcal{M}(\Omega)$ denote the set of finite measures defined on $\Omega$. The discretization operator $I:\mathcal{M}(\Omega)\to\mathbb{R}^{H\times H\times H}$ maps measures to discrete grid functions, defined for Dirac measures as:
(20) $I(\delta_{\mathbf{X}})=\sum_{\gamma\in\Gamma}F_{\gamma}\delta_{\mathbf{X}_{\gamma}},$
where $\Gamma$ is the interpolation basis index set from Section 3.1, $F_{\gamma}$ are the corresponding coefficients in (16) or (19), and $\mathbf{X}_{\gamma}$ denote grid positions.
Through linearity and Kantorovich-Rubinstein duality characterizing the Wasserstein distance $\mathcal{W}_{1}$, the operator $I$ extends uniquely via metric completion to all measures $\rho\in\mathcal{W}_{1}(\Omega)$ with finite mass. Namely, let $\rho_{k}\xrightarrow{\mathcal{W}_{1}}\rho$, where $\rho_{k}$ are discrete point measure approximations of $\rho$, we have $I(\rho)=\lim_{k\to\infty}I(\rho_{k})$.
Similarly, we define Grid-to-Particle interpolation.
@@
(21) $(I^{*}f)(\mathbf{X})=\sum_{\gamma\in\Gamma}F_{\gamma}f(\mathbf{X}_{\gamma}),$

### v3

@@ Abstract
In this paper, we develop a novel numerical framework, namely the stochastic interacting particle-field method with particle-in-cell acceleration (SIPF-PIC), for the efficient simulation of the three-dimensional (3D) parabolic-parabolic Keller-Segel (KS) systems. The SIPF-PIC method integrates Lagrangian particle dynamics with spectral field solvers by leveraging localized particle-grid interpolations and fast Fourier transform (FFT) techniques. For $P$ particles and $H$ Fourier modes per spatial dimension, the SIPF-PIC method achieves a computational complexity of $O(P+H^{3}\log H)$ per time step, a significant improvement over the original SIPF method (proposed in [SIPF1]), which has a computational complexity of $O(PH^{3})$, while preserving numerical accuracy. Moreover, we carry out a rigorous error analysis for the proposed method and establish rigorous corresponding error estimates. Finally, we present numerical experiments to validate the convergence order and demonstrate the computational efficiency of the SIPF-PIC method. To the best of our knowledge, this work represents the first numerical approach that can resolve complex 3D blowup dynamics beyond single-point collapse, including the formation of ring-shaped singular structures.
@@ 1 Introduction
Due to the nonlinear and potentially singular behavior of KS systems, especially in the presence of blow-up phenomena [nagai1995blow, herrero1998self], numerical methods have become essential for analyzing their solutions. Mesh-based numerical methods are among the most widely used approaches for solving the KS system. Chertock et al. [pos_preserving] proposed a high-order hybrid finite volume and finite difference scheme with positivity-preserving properties for two-dimensional (2D) problems. Shen et al. [shen2020unconditionally] proposed an energy dissipation and bound-preserving scheme that is not restricted to specific spatial discretizations. Chen et al. [chen2022error] developed a fully-discrete finite element scheme for the 2D parabolic-elliptic case, and demonstrated that the scheme exhibits finite-time blow-up under assumptions analogous to those of the continuous case. Other works include a positivity-preserving and asymptotic-preserving semi-discrete scheme [LiuJian-Guo2018Paap], and a saturation concentration setting [hillen2001global] which prevents blow-up and yields spiky solutions. Beyond these 2D cases, Gnanasekaran et al. [3Dblowup] simulated a 3D blow-up solution in a cube with a finite difference scheme in a short time.
@@
Existing Lagrangian frameworks for parabolic-elliptic cases cannot be directly extended to fully parabolic systems, as the chemo-attractant concentration no longer achieves rapid local equilibration and the drift terms depend on historical particle positions [chen2022mean, fournier2023particle]. Stevens [stevens2000] developed an $N$-particle system for the fully parabolic case, representing the chemo-attractant concentration with particles subject to probabilistic creation and annihilation. However, this method also suffers from the accumulation of per-step complexity over time. To address this limitation, Wang et al. [SIPF1] developed a stochastic interacting particle-field (SIPF) method that computes the chemo-attractant concentration through spectral decomposition and maintains only the current-step particle positions, which ensures that the per-step complexity remains constant over time.
@@
However, we have to point out that in the SIPF method [SIPF1], each particle of $\rho$ interacts with every Fourier mode of $c$. This requires the computation of the trigonometric sum $\frac{1}{P}\sum_{p=1}^{P}e^{-\frac{2\pi i\mathbf{q}}{L}\cdot\widehat{\mathbf{X}}_{p}}$ for all wavenumbers $\mathbf{q}$, where $\widehat{\mathbf{X}}_{p}$ for $1\leq p\leq P$ are the positions of the particles (see Algorithm 1). For 3D systems with $H$ Fourier modes per dimension, this results in an $O(PH^{3})$ computational complexity per timestep. Such a computational bottleneck severely restricts the applicability of SIPF for large-scale 3D KS systems, particularly in scenarios requiring high-resolution and long-time simulations.
@@
In this paper, to reduce the complexity, we propose the SIPF-PIC method (see Algorithm 4) that integrates localized particle-grid operations with spectral field solvers accelerated by the fast Fourier transform (FFT). This hybrid method confines particle interactions to adjacent grid nodes while resolving global frequency-domain couplings via FFT, yielding a computational complexity of $O\left(P+H^{3}\log H\right)$ per timestep, as presented in Theorem 3.7. The bidirectional interaction mechanism operates in each time step through two stages: particle-to-grid projection (see Algorithm 2) transfers particle-derived data to the computational grid, while grid-to-particle interpolation (see Algorithm 3) distributes field solutions from the grid back to individual particles; details are provided in Section 3. The particle-to-grid projection in the SIPF-PIC method, which numerically realizes the trigonometric summation that is also known as the unequally spaced fast Fourier transform (USFFT) [USFFT], has been employed in Fourier-basis particle simulations [USFFTS]. The grid-to-particle interpolation is employed in classical particle-in-cell methods [alma99, TskhakayaD.2007TPM] and does not rely on spectral decompositions. Our algorithm reduces the particle-grid interaction complexity from $O\left(PH^{3}\right)$ to $O\left(P+H^{3}\log H\right)$, thus enabling the simulation of 3D KS systems with significantly increased particle numbers $(P>10^{6})$ and finer grids $(H=256)$, and resulting in improved numerical accuracy.
@@
To complement the theoretical analysis above, we also present numerical experiments to verify the convergence rate, computational complexity, and consistency with respect to $H$. We verify the critical mass $8\pi$ and the evolution of the second moment in a standard two-dimensional parabolic-elliptic system (1), where the dimension $d=2$ and the parameters $(k,\epsilon,\mu,\chi)=(0,0,1,1)$. For 3D parabolic-parabolic KS systems with $k,\epsilon>0$, the critical mass lacks a precise characterization due to its dependence on the initial density distribution $\rho(\mathbf{x},0)$. We thus validate the accuracy of the SIPF-PIC method under radially symmetric initial data, where the reference solution can be obtained by reducing the system to 1D and running a finite difference method on a very fine mesh.
@@
Moreover, the SIPF-PIC method successfully reproduces complex blow-up patterns beyond the classical single-point radial type blow-up. To the best of our knowledge, this work presents the first numerical approach capable of resolving complex 3D blow-up dynamics beyond the single-point collapse regime, marking a key advance in the numerical simulation of such singular phenomena. We first investigate a non-radially symmetric configuration consisting of four uniform density balls located at the vertices of a tetrahedron, and then accurately resolve the ring-shaped blow-up pattern proposed in a recent paper [ring_ref], where mass concentrates on an evolving ring $x_{1}^{2}+x_{2}^{2}=r(t)^{2}$ with a time-dependent radius $r(t)$. In contrast, the original SIPF method, with parameter scalings ($P=10^{4},H=24$), lacks sufficient resolution to capture the ring-shaped blow-up pattern. Specifically, the SIPF-PIC method successfully captures the line-density blow-up, which is characterized by mass concentration along evolving spatial structures rather than isolated points, and further captures its instability under symmetry-breaking perturbations [collapsing]. These complex collapse mechanisms deviate fundamentally from classical delta-type blow-up dynamics, and the SIPF-PIC method, by leveraging large particle numbers and high-resolution grids, reliably captures their evolutionary processes in simulations. These results validate the performance of the SIPF-PIC method in addressing challenging 3D blow-up problems that were previously intractable by using existing methods.
@@
The rest of the paper is organized as follows. Section 2 presents theoretical finite-time blow-up results for the KS system (1), and introduces the SIPF method, a hybrid particle-field method derived through temporal discretization of the system. In Section 3, we formalize the interpolation operators governing bidirectional particle-grid coupling, including particle-to-grid discretization and grid-to-particle interpolation, followed by the detailed formulation of the SIPF-PIC method. Subsection 4 rigorously establishes a comprehensive error analysis framework that quantifies the approximation errors from various sources in the SIPF-PIC method. Section 5 validates the method through numerical experiments, including convergence tests, performance benchmarks, and the visualization of 3D blow-up patterns. Finally, a summary of our findings and future research directions is given in Section 6.
@@ 2.1 Analytical results of KS system
(3) $\frac{\mathrm{d}}{\mathrm{d}t}m_{2}(t)=4M_{0}\left(1-\frac{M_{0}}{8\pi}\right),$
@@ 2.2 SIPF Method
Throughout this section, the system is restricted to a domain $\Omega=\left[-\frac{L}{2},\frac{L}{2}\right]^{3}$ with periodic boundary conditions. In the SIPF method for solving the parabolic-parabolic KS system (2), the chemical concentration $\hat{c}(\mathbf{x},n\tau)$ is represented in frequency space using a Fourier series expansion, while the particle density $\hat{\rho}(\mathbf{x},n\tau)$ is represented in physical space by some moving particles. The temporal domain $[0,T]$ is partitioned into $N_{T}=\left\lceil\frac{T}{\tau}\right\rceil$ equal intervals of size $\tau$, where the ceiling function $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$.
@@
(5) $\hat{c}^{(n)}(\mathbf{x}):=\hat{c}(\mathbf{x},n\tau)=\sum_{\mathbf{q}\in U}\hat{\alpha}_{\mathbf{q}}^{(n)}e^{i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}},$
@@
where $\delta(\cdot)$ is the Dirac delta function, $M_{0}=\int_{\Omega}\rho^{(0)}(\mathbf{x})\,\mathrm{d}\mathbf{x}$ is the initial mass, and $(\widehat{\mathbf{X}}_{p}^{(n)})_{1\leq p\leq P}$ are positions of particles at time $t_{n}$.
@@
(9) $\hat{\beta}_{\mathbf{q}}^{(n-1)}=\frac{1}{L^{3}}\int_{\Omega}\hat{\rho}^{(n-1)}(\mathbf{x})e^{-i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}\,\mathrm{d}\mathbf{x}=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}e^{-\frac{2\pi i}{L}\mathbf{q}\cdot\widehat{\mathbf{X}}_{p}^{(n-1)}}.$
@@
(11) $\mathrm{d}\mathbf{X}=\chi\nabla_{\mathbf{X}}c(\mathbf{X},t)\mathrm{d}t+\mathrm{d}\mathbf{W}_{t}.$
@@
where $\mathbf{W}_{p}^{(n)}\sim\mathcal{N}(0,\,2\mu\tau I_{3})$ are independent 3D Wiener increments, $I_{3}$ is the $3\times 3$ identity matrix, and the drift $\nabla\hat{c}^{(n-1)}(\widehat{\mathbf{X}}_{p}^{(n-1)})$ is computed by
@@ 3.1 Particle-in-cell interpolation framework
The drawback of the original SIPF method (Algorithm 1) is the particle-frequency interaction with $O(PH^{3})$ computational complexity. Our new framework is designed to break through this computational bottleneck.
@@
Consider a smooth target function $f:\mathbb{R}^{3}\to\mathbb{R}$; $\mathbf{X}\in\Omega$ is an arbitrary coordinate. A conventional particle-in-cell (PIC) method [alma99] represents $f(\mathbf{X})$ as a combination of the function values at grid points:
(14) $f(\mathbf{X})\approx\sum_{\mathbf{q}\in U}R\left(\mathbf{X},\mathbf{q}\frac{L}{H}\right)f\left(\mathbf{q}\frac{L}{H}\right),$
where the particle-grid interaction kernel $R\left(\mathbf{X},\mathbf{q}\frac{L}{H}\right)$ satisfies $\sum_{\mathbf{q}\in U}R\left(\mathbf{X},\mathbf{q}\frac{L}{H}\right)=1$ for every $\mathbf{X}\in\Omega$. Furthermore, the kernel $R(\mathbf{X},\cdot)$ is locally supported, meaning that for a given $\mathbf{X}\in\Omega$, only a few nearby grid points contribute to the approximation.
For example, in a trilinear interpolation with a second-order truncation error, the kernel is represented as
(15) $R\left(\mathbf{X},\mathbf{q}\frac{L}{H}\right)=\prod_{1\leq j\leq 3}\max\left(1-\frac{H}{L}\left|X_{j}-q_{j}\frac{L}{H}\right|,0\right).$
@@
In particular, $R(\mathbf{X},\cdot)$ is nonzero only for the eight vertices of the grid cell containing $\mathbf{X}$. We denote the second-order basis $\Gamma_{2}=\{0,1\}^{3}$ and the relative grid positions $\mathbf{X}_{{\gamma}}=\left(\left\lfloor\frac{H}{L}\mathbf{X}\right\rfloor+{\gamma}\right)\times\frac{L}{H}$, where the notation $\lfloor\cdot\rfloor$ denotes the floor function applied component-wise. The interpolation (14) can be compactly expressed as
(16) $f(\mathbf{X})\approx\sum_{\boldsymbol{\gamma}\in\Gamma_{2}}F_{{\gamma}}^{(2)}(\boldsymbol{\lambda})f(\mathbf{X}_{{\gamma}}),$
@@
where the weights $F_{j_{1}j_{2}j_{3}}^{(2)}(\boldsymbol{\lambda})=\prod_{s=1}^{3}\left(1-j_{s}+(2j_{s}-1)\lambda_{s}\right)$ for $(j_{1},j_{2},j_{3})\in\Gamma_{2}$, and $\boldsymbol{\lambda}=\boldsymbol{\lambda}(\mathbf{X}):=\frac{H}{L}\mathbf{X}-\left\lfloor\frac{H}{L}\mathbf{X}\right\rfloor\in[0,1)^{3}$ is the relative location of $\mathbf{X}$ in the grid cell containing $\mathbf{X}$. For points near the boundary under periodic boundary conditions, the indices are wrapped around the grid edges.
To achieve higher precision, we also introduce the fourth-order basis
$\Gamma_{4}=\Gamma_{2}\cup\{-1,2\}\times\{0,1\}\times\{0,1\}\cup\{0,1\}\times\{-1,2\}\times\{0,1\}\cup\{0,1\}\times\{0,1\}\times\{-1,2\},$
which enables interpolation [tricubic] exhibiting a fourth-order truncation error, as shown in Figure 1.
Figure 1: Extended vertices for fourth-order.
The complete fourth-order interpolation formula becomes as follows:
(17) $f(\mathbf{X})\approx\sum_{\gamma\in\Gamma_{4}}F_{\gamma}^{(4)}(\boldsymbol{\lambda})f(\mathbf{X}_{\gamma}),$
where the weights $F_{\gamma}^{(4)}(\boldsymbol{\lambda})$ are shown below, exhibiting cubic symmetry:
(18) $F_{2,j_{2},j_{3}}^{(4)}(\boldsymbol{\lambda})=-\frac{(1+\lambda_{1})(1-\lambda_{1})\lambda_{1}\left(1-j_{2}+(2j_{2}-1)\lambda_{2}\right)\left(1-j_{3}+(2j_{3}-1)\lambda_{3}\right)}{6},$
$F_{-1,j_{2},j_{3}}^{(4)}(\boldsymbol{\lambda})=-\frac{(2-\lambda_{1})\lambda_{1}(1-\lambda_{1})\left(1-j_{2}+(2j_{2}-1)\lambda_{2}\right)\left(1-j_{3}+(2j_{3}-1)\lambda_{3}\right)}{6},$
$\cdots$
$F_{j_{1}j_{2}j_{3}}^{(4)}(\boldsymbol{\lambda})=\prod_{s=1}^{3}\left(1-j_{s}+(2j_{s}-1)\lambda_{s}\right)\left(1+\sum_{k=1}^{3}\frac{\lambda_{k}(1-\lambda_{k})}{2}\right),(j_{1},j_{2},j_{3})\in\{0,1\}^{3}.$
@@
[…]
(19) $I(\sum_{p=1}^{P}w_{p}\delta(\mathbf{x}-\widehat{\mathbf{X}}_{p}))=\sum_{p=1}^{P}w_{p}\sum_{\gamma\in\Gamma}F_{p,\gamma}\delta(\mathbf{x}-\widehat{\mathbf{X}}_{p,\gamma}),$
where $\Gamma$ is the interpolation basis index set from Section 3.1, $F_{p,\gamma}:=F_{\gamma}(\boldsymbol{\lambda}(\mathbf{X}_{p}))$ and $\widehat{\mathbf{X}}_{p,\gamma}:=\left(\left\lfloor\frac{H}{L}\mathbf{X}_{p}\right\rfloor+{\gamma}\right)\times\frac{L}{H}$ are the corresponding coefficients and grid points in (16) or (17) for $\widehat{\mathbf{X}}_{p}$.
Similarly, we define grid-to-particle interpolation as follows.
@@
(20) $(I^{*}f)(\mathbf{X})=\sum_{\gamma\in\Gamma}F_{\gamma}(\boldsymbol{\lambda}(\mathbf{X}))f(\mathbf{X}_{\gamma}),$